Inhaltsverzeichnis

[Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble] [Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble] [Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble], Sail; [Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble] [Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble] [Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble][Pleaseinsertintopreamble], Das Segel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Михаил Юрьевич Лермонтов, Sail; Михаил Юрьевич Лермонтов, Das Segel  2/2
Михаил Юрьевич Лермонтов,  Михаил Юрьевич Лермонтов,
Sail  Das Segel
tradukita de A. Z. Foreman  tradukita de Heinrich Greif

A sail drifts white and on its own Amid the light blue ocean haze. What does it seek in distant country? What made it leave its native bays?

Wo Meer und Himmel sich vereinen, Erglänzt ein Segel, weiß und weit - Was trieb es aus dem Land der Seinen? Was sucht es in der Einsamkeit?

Great billows play. High winds are whistling Down at the bending, creaking mast Oh! This one seeks no happy ending And does not flee a happy past.

Es pfeift der Wind. Die Wellen drohen. Es knarrt der Mast. Das Segel schwebt Nicht vor dem Glück ist es geflohen. Es ist nicht Glück, wonach es strebt.

Beneath, a brighter stream than azure. Above, the golden sunray flows Yet this one, restive...quests for tempests As if in tempests were repose.

Strahlt auch in Gold der Himmelsbogen, Und glänzt auch noch so blau das Meer - Das Segel lechzt nach Sturm und Wogen, Als ob in Stürmen Ruhe wär.

Traduko de la Rusa poemo “Парус” de Михаил Юрьевич Лермонтов (∗1814-10-16 – †1841-07-

27) en la Anglan de A. Z. Foreman.

Arg-545-2269 (2015-01-28 19:03:20)

I, Manfred Retzlaff, found this poem-translation in http: // poemsintranslation. blogspot. e/ 2009/ 10/ lermontov-sail-from-russian. html .

Traduko de la Rusa poemo “Парус” de Михаил Юрьевич Лермонтов (∗1814-10-16 – †1841-07-27) en la Germanan de Heinrich Greif (∗1907-03-11 – †1946-07-16).

Arg-545-1944 (2015-01-24 19:29:11)

Pri Heinrich Greif vidu la retejon http: // de. wikipedia. org/ wiki/ Heinrich_Greif . La traduko troviĝas en http: // langadvice. blogspot. ru/ 2012/ 10/ blog-post. html .

http://www.poezio.net/version?poem-id=545&version-id=2269,1944